### v1

@@ 5.3 Ablation Study
Diffusion vs. Flow. We show that employing Flow Matching for the generative model is essential. To prove that, we train LayoutFlow using a diffusion loss instead of the proposed flow matching loss. For sampling from diffusion-based models, we test two methods, i.e., DDPM and DDIM. As shown in Tab. 4 neither of them is able to reach the same performance as LayoutFlow and both seem to be particularly lacking in terms of overlap. Interestingly, DDIM outperforms DDPM by a large margin which might be related to DDIM sampling based on an ODE instead of an SDE.
Regularization Loss. We validate the efficacy of our regularization loss proposed in Eq. 6 and the choice of $\lambda$ in Tab. 4. It can be clearly observed that a stronger regularization using the $L1$-loss provides better alignment but increases the FID. For that reason, we decided to consider the trade-off across all metrics and chose $\lambda=0.2$ for LayoutFlow.
Table 3: Ablation study on the choice of training approaches using the same architecture.
@@
Table 4: Ablation study on different loss functions.
@@
Initial Distribution. Since flow-based models allow for arbitrary initial distributions, we explore two alternatives. First, we try a uniform distribution to ensure all initial samples are placed on the canvas. In addition, as Analog Bits have a smaller sample space, we tested using a Gaussian for the geometrical elements and a uniform distribution for the Analog Bits. Overall, the results in Tab. 6 indicate that a Gaussian initial distribution provides the best overall performance measured by the FID score, whereas a uniform distribution provides stronger results in terms of geometrical metrics, such as alignment and overlap.
Table 5: Ablation study on different initial distributions.
@@
Table 6: Ablation study on different training trajectories.
@@
Cosine 2.48 0.152 0.480 0.565
Training Trajectories. As long as the conditional vector field fulfills the criteria presented in Sec. 3, it can be freely chosen for training. In addition to the linear training trajectory proposed in [32, 44], we also explore the sine/cosine interpolation introduced in [1] and a cosine-based interpolation, as shown in Tab. 6. The choice of training trajectories only slightly affects the performance.
@@ 6 Conclusion
In this paper, we explored the application of Flow Matching for layout generation and as a result proposed LayoutFlow, a flow-based model that is able to handle various layout generation tasks. Our model is able to significantly speed up inference compared to other diffusion-based models while providing state-of-the-art performance. Even though we tested a diverse set of design options, there still remains room for further exploration into applying Flow Matching to layout generation, for example, different training trajectories, conditioning methods, or initial distributions. Furthermore, LayoutFlow might also be extended to generate layouts considering the content in the elements, similar to related work [46, 19]. Overall, we demonstrate that Flow Matching provides a highly flexible and powerful tool for layout generation that offers a natural geometrical interpretation.
@@
In this supplementary material, we provide additional explanations and evaluations to supplement the contents of the main paper. Specifically, we present further discussions on training trajectories, conditioning, and the limitations of our method, as well as additional quantitative and qualitative results.
@@ Appendix 0.A Training Trajectories
The conditional flow, or training trajectory, $\phi_{t}(\mathbf{x})=\mathbf{x}_{t}$ can be chosen freely as long as the conditional vector field follows the continuity equation and $\phi_{0}(\mathbf{x})=\mathbf{x}_{0}$ and $\phi_{1}(\mathbf{x})=\mathbf{x}_{1}$. While LayoutFlow is trained using a linear trajectory following [32, 44], we also tested the sine/cosine interpolation introduced in [1] and a cosine-based interpolation. More details are shown in Tab. 7. Our main motivation for the cosine-based interpolation was to preserve the direction of the conditional vector field but apply a different time schedule to the trajectory. In the linear scenario, the conditional vector field remains independent of the time and always points towards $x_{1}$. On the other hand, the sine/cosine trajectory changes its direction over time, only gradually leading to $x_{1}$ in a circular trajectory. The cosine-based interpolation maintains the direction of the linear trajectory but weighs it depending on the time. At an early time of the trajectory, the magnitude of the direction is large, whereas towards the end the derivative starts to slow down. In analogy to the time importance sampling in diffusion models [28], our hypothesis was that this might help focus on the later stages of the flow process during training where more detailed features become important. However, based on the paper’s ablation results, the cosine-based interpolation does not seem to affect the performance. Nonetheless, this scheduling property of conditional fields might be further explored with other functions of the general form:
@@
while still fulfilling the conditions mentioned above. In the cosine case of $\kappa(t)=\cos(\frac{pi}{2}t)$, the schedule might have been too close to the linear schedule to make a noticeable change and different choices might provide better results.
Table 7: Overview of different training trajectories and conditional vector fields. Note that the conditional vector field is the time-derivative of the training trajectory
@@
Cosine $(1-\sin{(\frac{\pi}{2}t}))\mathbf{x}_{0}+\sin{(\frac{\pi}{2}t)}\mathbf{x}_{1}$ $\cos{(\frac{\pi}{2}t)}(\mathbf{x}_{1}-\mathbf{x}_{0})$
Appendix 0.B Quality Speed Trade-off
In Fig. 8, we compare the number of function evaluations of our model to the performance on the RICO dataset, corresponding to the number of steps taken to solve the ODE. While only 5 function evaluations result in a bad score of an FID of roughly 50, just three more evaluations can bring down the FID to 13. The improvement observed by increasing the number of steps plateaus at around 40. In comparison, diffusion models typically require at least 100 steps to perform well. Overall, LayoutFlow offers a strong quality-speed trade-off, where slightly reducing the number of steps, e.g. from 30 to 20, only has a small impact on performance.
@@ Appendix 0.B Quality Speed Trade-off
Appendix 0.C Conditioning Analysis
In general, there are several ways to introduce conditions to a generative task. For layout generation, given conditions can differ significantly depending on the task. For LayoutFlow, we employ a masking strategy during training that indicates which part of the input serves as a condition, similar to  [5, 20, 28]. Intuitively, the masking mechanism can be considered a switch that allows training multiple flow-based task-specific models with a single shared architecture in just one model. Alternatively, it is possible only to train an unconditional model and impose the conditions during sampling through a guidance strategy as done in  [22, 48]. This strategy only requires training the model on the unconditional task and inserting the condition during inference, effectively altering the trajectory.
To condition the trajectory during inference, we employ a mask $\mathbf{m}$ that is zero for the condition dimension, i.e. all ones in the case of unconditional generation. Essentially, we use the mask only to update the direction of the condition dimension to $\mathbf{u^{\prime}}$ and can be described as
@@ Appendix 0.C Conditioning Analysis
Appendix 0.D Results Using Different Data Split
Due to the different data splits used by previous methods, comparing the numbers reported by other papers has become difficult. The FID values depend on a separately trained network, making comparing even more difficult as the same weights and network are required. Our experiments used the same settings as in [48, 23]. As a result, we had to retrain LayoutDM [22] and DLT [28] with that specific data split. To ensure that LayoutFlow works robustly across different dataset settings, we also trained a model on the data split used in LayoutDM [4] which includes up to 25 elements per layout as opposed to only 20 elements in the other setting. Additionally, this allows us to compare with a concurrent work called LACE [5], which uses a continuous diffusion model.
@@ Appendix 0.D Results Using Different Data Split
Appendix 0.E Completion Task
The completion task describes generating a complete layout given an incomplete layout. In this scenario, the given incomplete layout acts as a strong condition. There are different ways to define the completion task, e.g. Inoue et al. assume that up to 20% of the layout is already given. While this scenario is closer to unconditional generation as it still offers a high degree of flexibility, we also looked into completing almost finished layouts containing more than 80% of their elements. To show that LayoutFlow can handle either scenario, we train the model with the respective conditional mask and present the results in Tab. 10 compared to LayoutDM and show the qualitative results in Tab. 16.
@@ Appendix 0.E Completion Task
Appendix 0.F Limitations and Broader Impact
@@
Our proposed model, LayoutFlow, has shown remarkable results. However, some areas can still be improved in the future. The Alignment remains one of the largest challenges, especially for models working in the continuous data space. While LayoutFlow utilizes an L1 regularization loss to improve the alignment, the issue is not fully resolved. For the future, it is crucial to find a better loss function that can reflect better the perceptual error between layouts rather than relying too heavily on metrics that do not necessarily correlate with important design aspects such as alignment.
Appendix 0.G More Qualitative Results
@@
DLT LayoutDM LayoutDiffusion LayoutFlow

### v2

@@ 5.3 Ablation Study
Diffusion vs. Flow. We show that employing Flow Matching as the generative model is essential by also training the model used for LayoutFlow and only substituting Flow Matching with diffusion. We test two methods for sampling from the diffusion-based model, i.e., DDPM and DDIM. As shown in Tab. 4, neither approach is able to reach the same performance as LayoutFlow, and both seem to be particularly lacking in terms of Overlap. Interestingly, DDIM outperforms DDPM by a large margin, which might be related to DDIM sampling based on an ODE instead of an SDE.
Regularization Loss. We validate the efficacy of our regularization loss proposed in Eq. 6 and the choice of $\lambda$ in Tab. 4. It can be clearly observed that a stronger regularization using the $L1$-loss provides better alignment but increases the FID. Therefore, we considered the trade-off across all metrics and chose $\lambda=0.2$ for LayoutFlow.
Table 3: Ablation study on the choice of training approaches using the same architecture
@@
Table 4: Ablation study on different loss functions
@@
Initial Distribution. Since flow-based models allow for arbitrary initial distributions, we explore two alternatives. First, we try a uniform distribution to ensure all initial samples are placed on the canvas. In addition, as Analog Bits have a smaller sample space, we tested using a Gaussian for the geometrical elements and a uniform distribution for the Analog Bits. Altogether, the results in Tab. 6 indicate that a Gaussian initial distribution provides the best overall performance measured by the FID score, whereas a uniform distribution provides stronger results in terms of geometrical metrics, such as Alignment and Overlap.
Table 5: Ablation study on different initial distributions
@@
Table 6: Ablation study on different training trajectories
@@
Sine 2.48 0.152 0.480 0.565
Training Trajectories. As long as the conditional vector field fulfills the criteria presented in Sec. 3, it can be freely chosen for training. In addition to the linear training trajectory proposed in [32, 44], we also explore the sine/cosine interpolation introduced in [1] and a sine-based interpolation, as shown in Tab. 6. The choice of training trajectories only slightly affects the performance.
@@ 6 Conclusion
In this paper, we explored the application of Flow Matching for layout generation and, as a result, proposed LayoutFlow, a flow-based model that is able to handle various layout generation tasks. Our model is able to significantly speed up inference compared to other diffusion-based models while providing state-of-the-art performance. Even though we tested a diverse set of design options, there still remains room for further exploration into applying Flow Matching to layout generation, for example, different training trajectories, conditioning methods, or initial distributions. Furthermore, LayoutFlow might also be extended to generate layouts considering the content in the elements, similar to related work [46, 19]. Overall, we demonstrate that Flow Matching provides a highly flexible and powerful tool for layout generation that offers a natural geometrical interpretation.
Acknowledgments
This work was supported by JSPS KAKENHI Grant Number JP23K28139.
@@ Acknowledgments
In this document, we provide additional explanations and evaluations to supplement the contents of the main paper. Specifically, we present further discussions on training trajectories, conditioning, and the limitations of our method, as well as additional quantitative and qualitative results.
@@ Appendix 0.A Training Trajectories
The conditional flow, or training trajectory, $\phi_{t}(\mathbf{x})=\mathbf{x}_{t}$ can be chosen freely as long as the conditional vector field follows the continuity equation and $\phi_{0}(\mathbf{x})=\mathbf{x}_{0}$ and $\phi_{1}(\mathbf{x})=\mathbf{x}_{1}$. While LayoutFlow is trained using a linear trajectory following [32, 44], we also tested the sine/cosine interpolation introduced in [1] and a sine-based interpolation. More details are shown in Tab. 7. Our main motivation for the sine-based interpolation was to preserve the direction of the conditional vector field but apply a different time schedule to the trajectory. In the linear scenario, the conditional vector field remains independent of the time and always points towards $x_{1}$. On the other hand, the sine/cosine trajectory changes its direction over time, only gradually leading to $x_{1}$ in a circular trajectory. The sine-based interpolation maintains the direction of the linear trajectory but weighs it depending on the time. At an early time of the trajectory, the magnitude of the direction is large, whereas towards the end, the derivative starts to slow down. In analogy to time importance sampling in diffusion models [28], we hypothesized that this might help focus on the later stages of the flow process during training, where more detailed features become important. However, based on the paper’s ablation results, the cosine-based interpolation does not seem to affect the performance. Nonetheless, this scheduling property of conditional fields might be further explored with other functions of the general form:
@@
while still fulfilling the conditions mentioned above. In the cosine case of $\kappa(t)=\cos(\frac{pi}{2}t)$, the schedule might have been too close to the linear schedule to make a noticeable change, and different choices might provide better results.
Table 7: Overview of different training trajectories and conditional vector fields. Note that the conditional vector field is the time-derivative of the training trajectory.
@@
Sine $(1-\sin{(\frac{\pi}{2}t}))\mathbf{x}_{0}+\sin{(\frac{\pi}{2}t)}\mathbf{x}_{1}$ $\cos{(\frac{\pi}{2}t)}(\mathbf{x}_{1}-\mathbf{x}_{0})$
Appendix 0.B Inference Speed Analysis
Since previous layout generation approaches rely on different architectures, we provide some insights on how the architectural components, i.e., the model size measured by the numbers of trainable parameters and the number of tokens needed to represent a single layout, influence the inference speed in Table 8. In addition, we also report the number of steps each method requires to generate a layout that is tied to the choice of the generative model. Overall, it can be seen that the speed-up for LayoutFlow can be attributed to architectural improvements as well as changing from a diffusion model to a flow-base model due to the reduced number of generation steps required.
Table 8: Overview of different factors contributing to inference speed
Model Size Token per Layout Generation Steps Inference Speed
LayoutDiffusion 86M $>$100 160 1600.00ms
LayoutDM 12.4M 100 100 16.60ms
DLT 9.0M 20 100 3.50ms
LayoutFlow (ours) 12.7M 20 50 1.75ms
Appendix 0.C Quality Speed Trade-off
In Fig. 8, we compare the number of function evaluations of our model, corresponding to the number of steps taken to solve the ODE, to the performance on the RICO dataset. While only 5 function evaluations result in a bad score of an FID of roughly 50, just three more evaluations can bring down the FID to 13. The improvement observed by increasing the number of steps plateaus at around 40. In comparison, diffusion models typically require at least 100 steps to perform well. Overall, LayoutFlow offers a strong quality-speed trade-off, where slightly reducing the number of steps, e.g., from 30 to 20, only has a small impact on performance.
@@ Appendix 0.C Quality Speed Trade-off
Appendix 0.D Conditioning Analysis
There are several ways to introduce conditions to a generative task. For layout generation, given conditions can differ significantly depending on the task. For LayoutFlow, we employ a masking strategy during training that indicates which part of the input serves as a condition, similar to  [5, 20, 28]. Intuitively, the masking mechanism can be considered a switch that allows training multiple flow-based task-specific models with a single shared architecture in just one model. Alternatively, it is possible only to train an unconditional model and impose the conditions during sampling through a guidance strategy as done in  [22, 48]. This strategy only requires training the model on the unconditional task and inserting the condition during inference, effectively altering the trajectory.
To condition the trajectory during inference, we employ a mask $\mathbf{m}$ that is zero for the condition dimension, i.e., all ones in the case of unconditional generation. Essentially, we use the mask only to update the direction of the condition dimension to $\mathbf{u^{\prime}}$ and can be described as
@@ Appendix 0.D Conditioning Analysis
Appendix 0.E Results Using Different Data Split
Due to the different data splits used by previous methods, comparing the numbers reported by other papers has become difficult. The FID values depend on a separately trained network, making comparing even more difficult as the same weights and network are required. Our experiments used the same settings as in [48, 23]. As a result, we had to retrain LayoutDM [22] and DLT [28] with that specific data split. To ensure that LayoutFlow works robustly across different dataset settings, we also trained a model on the data split used in LayoutDM [4], which includes up to 25 elements per layout as opposed to only 20 elements in the other setting. Additionally, this allows us to compare with a concurrent work called LACE [5], which uses a continuous diffusion model.
@@ Appendix 0.E Results Using Different Data Split
Appendix 0.F Completion Task
The completion task describes generating a complete layout given an incomplete layout. In this scenario, the given incomplete layout acts as a strong condition. There are different ways to define the completion task, e.g., Inoue et al. assume that up to 20% of the layout is already given. While this scenario is closer to unconditional generation as it still offers a high degree of flexibility, we also looked into completing almost finished layouts containing more than 80% of their elements. To show that LayoutFlow can handle either scenario, we seperately trained models with the respective conditional mask and present the results in Tab. 11 compared to LayoutDM and show the qualitative results in Appendix 0.H.
@@ Appendix 0.F Completion Task
Appendix 0.G Limitations and Broader Impact
@@
Our proposed model, LayoutFlow, has shown remarkable results. However, some areas can still be improved in the future. Alignment remains one of the largest challenges, especially for models working in the continuous data space. While LayoutFlow utilizes an L1 regularization loss to improve the alignment, the issue is not fully resolved. For the future, it is crucial to find a better loss function that can reflect the perceptual error between layouts rather than relying too heavily on metrics that do not necessarily correlate with important design aspects such as alignment.
Appendix 0.H More Qualitative Results
@@
DLT LayoutDM LayoutFlow Ground Truth